OpenStudy (anonymous):

A babysitter charges 25% more than her normal hourly rate on Saturday night. Her normal hourly rate is $12.50. Which expression represents how much the babysitter charges on Saturday night?

OpenStudy (anonymous):

\[12.50\times 5/4\] \[12.50 \times 7/4\]

OpenStudy (anonymous):

\[12.50 \times 1/4\] \[12.50\times 3/4\]

OpenStudy (sydthekid913):

It is C. because 1/4 is equivalent to 25%

OpenStudy (sydthekid913):

and to find the percentage of a number we multiply it by the percent we are trying to find

OpenStudy (anonymous):

C

OpenStudy (sydthekid913):

her hourly rate on a Saturday night would be 12.50 * .25 = 3.125 + 12.50 = 15.625